OpenStudy (anonymous):

Given PR=RT, use the properties of equality for segment length to show that RT=PQ+QR.

OpenStudy (aum):

Can you draw the diagram using the Draw button below the reply box? Or you can post a screenshot of the question.

OpenStudy (anonymous):

This is addition property

OpenStudy (aum):

PR = RT (given) PR = PQ + QR Therefore, RT = PQ + QR

OpenStudy (anonymous):

Can you help me with #31?

OpenStudy (aum):

3x - 4 = 2x + 2 3x - 4 + 4 = 2x + 2 + 4 (addition property of equality) 3x = 2x + 6 (simplification) 3x - 2x = 2x + 6 - 2x (subtraction property of equality) x = 6 (simplification)

OpenStudy (aum):

If you wish you can skip the simplification step and write it as follows: 3x - 4 = 2x + 2 3x = 2x + 6 (addition property of equality) x = 6 (subtraction property of equality)

OpenStudy (anonymous):

For #31 question A Do you think I did it right or wrong?